OpenStudy (anonymous):

Solve the triangle. B = 73°, b = 15, c = 10

OpenStudy (anonymous):

C= 39.61 degrees A= 67.39 degrees a= 14.48

OpenStudy (anonymous):

Thank you!! @moe_ha

OpenStudy (anonymous):

can you help me with another one?...

OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. C = 67°, a = 21, c = 20

OpenStudy (anonymous):

You're welcome. wait i think theyre is another part to the first one.... Yeah I can help with the other one. hold on.

OpenStudy (anonymous):

can you explain what you did on the first one?

OpenStudy (anonymous):

yeah sure so on the first one i used the law of sines. so \[\frac{ \sin73 }{ 15 } =\frac{ sinC }{ 10 } \] then \[\frac{ 10\sin73 }{ 15 } \] which equals .637536504 then i did the arcsine of that which gave me C=39.61 then just add 73 and 39.61 and subtract that from 180 and that'll give you your A and that equals 67.39. to find the side a \[\frac{ \sin73 }{ 15 } \frac{ \sin67.39 }{ a }\] that'll give you \[\frac{ 15\sin67.39 }{ \sin73 }\] it'll equal 14.48 and that's how i did the first question

OpenStudy (anonymous):

This is for the first triangle:\[\frac{ \sin67 }{ 20 } \frac{ sinA }{ 21 } = \frac{ 21\sin67 }{ 20 }\] then use the arcsine to find A A should equal 75.13 degrees add 75.13 and 67 the subtract it from 180 it = 37.87 degrees and that is your B \[\frac{ \sin67 }{ 20 } \frac{ \sin37.87 }{ b } =\frac{ 20\sin37.87 }{ \sin67 } =13.34\] For the second triangle: subtract 75.13 from 180 so this will give you a new A = 104.87 degrees to find your new a you put it back into the law of sine formula \[\frac{ \sin67 }{ 20 } \frac{ \sin104.87 }{ b } = \frac{ 20\sin104.87 }{ \sin67 } = 21.00\] side a=21.00 to find your new B you add 67 and 140.87 which gives you 8.13 degrees \[\frac{ \sin67 }{ 20 } \frac{ \sin8.13 }{ b } = \frac{ 20\sin8.13 }{ \sin67 } = 3.07\] so your new side b = 3.07

OpenStudy (anonymous):

so for your first triangle A= 75.13 degrees B= 37.87 degrees b= 13.34 for the second triangle A= 104.87 degrees a= 21.00 B= 8.13 degrees b= 3.07 hoped I helped :)

OpenStudy (anonymous):

Thanks so much! @moe_ha

OpenStudy (anonymous):

you're welcome!! if you have anymore questions just ask me :) i'll try to answer them

OpenStudy (anonymous):

well there is one more but theres a picture and I can't post the picture.. its just to help.. can you give it a shot? @moe_ha

OpenStudy (anonymous):

shoot. I'll try it.

OpenStudy (anonymous):

Two weather tracking stations are on the equator 159 miles apart. A weather balloon is located on a bearing of N 38°E from the western station and on a bearing of N 14°E from the eastern station. How far is the balloon from the western station?

OpenStudy (anonymous):

oh god...word problem... i cannot do these. i completely suck at them I'm sorry.

OpenStudy (anonymous):

no worries! @moe_ha what about Solve the triangle. A = 32°, a = 19, b = 12 A. B = 19.6°, C = 148.4°, c ≈ 22.5 B. B = 19.6°, C = 128.4°, c ≈ 28.1 C. Cannot be solved D. B = 19.6°, C = 128.4°, c ≈ 16.9

OpenStudy (anonymous):

sorry for answering so late @ChandlerT0125 i was working on a paper. so the answer is b

OpenStudy (anonymous):

Thank you! @moe_ha

OpenStudy (anonymous):

You're welcome